OpenStudy (anonymous):

write polynomial function with given roots: 2i and - square root 2

OpenStudy (anonymous):

\[-\sqrt{2}\]

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Mertsj help plzz :/

OpenStudy (mertsj):

Now remember...if 2i is a root, then -2i is also a root. Furthermore is -sqrt2 is a root then +sqrt2 is also a root. Does that help?

OpenStudy (anonymous):

yes but how would i set it up i already have that part set

OpenStudy (mertsj):

Do you know that if x=3 is a root then x-3 is a factor?

OpenStudy (mertsj):

So if x = 2i is a root then x-2i is a factor. So your equation must be: \[(x-2i)(x+2i)(x-\sqrt{2})(x+\sqrt{2})=0\]

OpenStudy (mertsj):

Which in simpler form is : \[(x^2+4)(x^2-2)=0\]

OpenStudy (anonymous):

ok and if you dont do it in simpler form do i have to set it to zero

OpenStudy (mertsj):

Which is the same as: \[x^4+2x^2-8=0\]

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

Or you could write f(x)= because the directions say to write a polynomial function.

OpenStudy (anonymous):

ok and then i guess thats where i have to start working it out

OpenStudy (mertsj):

So If you just write the factors like I did and multiply them out you will get f(x)=x^4+2x^2-8 and that is your answer.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (mertsj):

yw